If the field strength is E = 9 V/m a distance of 1 m from the charge, what is the field strength E a distance of 3 m from the charge?

E =
Answer #1

E=kq/r^2

E=kq/r^2
at r=1
E=kq=9
at r=3
E=kq/r^2=9/3^3=9/9=1V/m